OpenStudy (anonymous):

sum of three consecutive odd integers is 207. What are the integers.

OpenStudy (campbell_st):

let the ist integer be x then next odd is x + 2 and the next is x + 4 then x + x +2 + x + 4 = 207 3x + 6 = 207 subtract 6 from both sides 3x = 201 I'll let you solve for x

OpenStudy (anonymous):

i dont know how to solve this.

OpenStudy (campbell_st):

divide both sides by 3

OpenStudy (anonymous):

can you show me how, then i will remember how do it next time?

OpenStudy (campbell_st):

if you have a computer then you have a calculator... now just do the calculation

OpenStudy (anonymous):

207 divided by 3 = 69, now how do i find out the second integer

OpenStudy (campbell_st):

its 201/3 = 67 the integers are 67, 69 and 71